OpenStudy (laddiusmaximus):

cos(pi+x)/cos((3pi/2)-x)=cotx prove

OpenStudy (anonymous):

[cos(pi)cosx - sin(pi)sinx]/[cos(3pi/2)cosx + sin(3pi/2)sinx] =(-cosx)/(-sinx) =cotx

OpenStudy (anonymous):

\[\cos(x+\pi)=-cos x\] \[\cos(3\pi/2-x)=-sin x\] \[-\cos x/-\sin x=\cot x\]